OpenStudy (anonymous):

3^2b^2c^-4 times 3^-1b^-5c^7 as a fraction

OpenStudy (anonymous):

\[\large3^2b^2c^{-4}\times 3^{-1}b^{-5}c^7\] \[=\huge 3^{2-1}b^{2-5}c^{-4+7}\]

OpenStudy (anonymous):

how would it look like as a fraction?

OpenStudy (anonymous):

you should get \[\large 3b^{-3}c^3\]

OpenStudy (anonymous):

if you have to write with positive exponents, note that \[b^{-n}=\frac{1}{b^n}\] and so \[b^{-3}=\frac{1}{b^3}\] making your answer \[\large \frac{3c^3}{b^3}\]